OpenStudy (accidentalaichan):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Part 3: Using your equation from part 2, rewrite this equation in standard form.

OpenStudy (anonymous):

start with \[y-y_1=m(x-x_1)\] with \[x_1=3,y_1=-2, m=2\] to get the "point slope" form

OpenStudy (accidentalaichan):

y - (-2) = 2(x - 3)?

OpenStudy (anonymous):

yes, but you probably want to write \[y+2=2(x-3)\] that is the "point slope" form to turn that into the "slope intercept" form, solve the above equation for \(y\)

OpenStudy (accidentalaichan):

y = mx + b y = 2/-3x + 2? Did I do it right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[y+2=2(x-3)\] steps always the same 1) remove the parentheses using the distributive law on the right \[y+2=2x-6\] then subtract \(2\) from both sides \[y=2x-8\]